Solve the system
6x-5y=-1
2y-2x=0

Answer by Seutip(231)

(1)6x-5y=-1
(2)2y-2x=0
(2)2y-2x=0 Find "y"
(2)2y=2x
(2)y=x
(1)6x-5y=-1
(1)6x-5x=-1 Substitute "x" to "y"
(1)x=-1
therefore,
y=-1
Since y=x


Hope that helps!
Pls do follow me on insta and twitter @seutip!

<3